OpenStudy (beth12345):

The gravitational field strength is also governed by an inverse square law. What is the gravitational field strength 200km above the Earth's surface, at the altitude of many manned space flights?

OpenStudy (beth12345):

The answer I'm given is 9.2N/kg but U don't know how to get it. I also have an equation sheet with the following equations... \[F _{g}=m \times g\]\[F=G \times ((m _{1} \times m _{2})/r ^{2})\]\[F _{fr}= \mu \times F _{N}\]\[F _{net}= m \times a\]\[p=m \times v\]\[\Delta p= F \times \Delta t\]\[F=k \times x\] but I also don't know what one to use, if someone could help me, that would be nice :)

OpenStudy (beth12345):

I* don't know how to get it

OpenStudy (apoorvk):

okay in the formula for gravitation force just add to 200km to the earth's radius of 6400kms... and calculate... although there is a more generic formula http://en.wikipedia.org/wiki/Gravity_of_Earth#Altitude

OpenStudy (beth12345):

so would I use the second equation?

OpenStudy (beth12345):

or the first?

OpenStudy (apoorvk):

you mean on the wiki page? ofcourse the first because you need to find it for above the surface, at an 'altitude'

OpenStudy (apoorvk):

in second equation if you need it for future questions, take rho nought and rho 1 as same, so the cancel out.

OpenStudy (beth12345):

so is the equation \[g _{h}=g _{0}\left[\begin{matrix}r e /\\ re + h\end{matrix}\right]^{2}\]?